# LAGUNA: LAnguage Guided UNsupervised Adaptation with structured spaces

### v2

@@ Abstract
Unsupervised domain adaptation remains a critical challenge in enabling the knowledge transfer of models across unseen domains. Existing methods struggle to balance the need for domain-invariant representations with preserving domain-specific features, which is often due to alignment approaches that impose the projection of samples with similar semantics close in the latent space despite their drastic domain differences. We introduce LAGUNA - LAnguage Guided UNsupervised Adaptation with structured spaces, a novel approach that shifts the focus from aligning representations in absolute coordinates to aligning the relative positioning of equivalent concepts in latent spaces. LAGUNA defines a domain-agnostic structure upon the semantic/geometric relationships between class labels in language space and guides adaptation, ensuring that the organization of samples in visual space reflects reference inter-class relationships while preserving domain-specific characteristics. Remarkably, LAGUNA surpasses previous works in $18$ different adaptation scenarios across four diverse image and video datasets with average accuracy improvements of $+3.32\%$ on DomainNet, $+5.75\%$ in GeoPlaces, $+4.77\%$ on GeoImnet, and $+1.94\%$ mean class accuracy improvement on EgoExo4D.
@@ 1 Introduction
Based on this observation, we introduce LAGUNA - LAnguage Guided UNsupervised Adaptation with structured spaces, a novel approach to domain adaptation which guides source and target spaces to develop semantic-geometric inter-relationships reflecting the structure of a reference space (Figure 1(bottom)). As shown in recent works, language can provide a semantic space agnostic to the nuances of visual observations enabling robust zero-shot generalization [54, 10] and supporting domain robustness [14, 66, 23, 45, 27], hence we choose language to build our reference structured space in LAGUNA. This assumes the availability for both source and target samples of natural language descriptions, which can be generated from available captioning models [32] or collected from the web at a fraction of the cost of typical human labeling procedures [27]. Specifically, LAGUNA employs a 3-stage approach to structurally align the source and target representation spaces while allowing each domain to preserve its typical patterns. In Stage 1, textual class labels are mapped to a domain-agnostic reference space representing their relationships at the conceptual level. In Stage 2, a language model is trained to map textual captions to the reference latent space in order to provide pseudo-labels for target samples Finally, in Stage 3, a cross-domain classifier is trained to encourage domain-specific representations to follow the reference structure.
Experiments demonstrate the superiority of LAGUNA over existing state-of-the-art methods across $18$ different domain adaptation scenarios sourced from four diverse image and video datasets, with gains of $+3.32\%$ on DomanNet [50], $+5.75\%$ on GeoPlaces [26], $+4.77\%$ on GeoImnet [26] and $+1.94\%$ mean per class accuracy on EgoExo4D [20]. We further report ablations to analyze the specific contributions of our design choices.
In sum, our main contributions are as follows: 1) we investigate the use of relative representations for UDA, showing its advantages with respect to the current absolute alignment paradigm; 2) we propose LAGUNA, a three-stage method which learns a cross-domain classifier where individual source and target spaces are distinct yet aligned to a reference structure derived from language; 3) through extensive ablations and comparisons with state-of-the-art models, we show the superiority of the proposed approach. Code available at: https://github.com/ADiko1997/LAGUNA.git
@@ 2.1 UDA in Computer Vision
Figure 2: The 3 stages of LAGUNA. In Stage 1, a domain-agnostic semantic structure in the form of a set of anchors $\mathcal{A}$ is defined. In Stage 2, a language model $\mathcal{G}$ is trained on source captions-label pairs to provide pseudo-labels for the target domain. The model is trained to be structurally aligned to anchors ${}^{*}\mathcal{A}$ with a structural loss $\mathcal{L}_{S}$ and to predict class labels with a cross-entropy loss $\mathcal{L}_{CE}$. In Stage 3, a cross-domain visual classifier is trained on labeled target and unlabeled source images. A visual encoder $\mathcal{V}$ is used to represent target and source samples as $g_{i}^{t}$ and $g_{i}^{s}$. The cosine similarities $\mathbf{r}^{g_{i}^{t}}$ between the target visual representation $g_{i}^{t}$ and all learnable target anchors $\mathcal{A}_{t}$ are bound to be similar to the cosine similarities $\mathbf{r}^{z_{i}^{t}}$ between the textual representation of the target caption $z_{i}^{t}$ predicted through the language model $\mathcal{G}$ and the reference domain-agnostic anchors ${}^{*}\mathcal{A}$ by means of the structural loss $\mathcal{L}_{S}$. A similar processing is applied to the visual representation of source sample $g_{i}^{s}$, the textual representation of the ground truth class ${}^{*}A[y_{i}^{S}]$ and the learnable source anchors $\mathcal{A}_{s}$. A regularization loss $\mathcal{L}_{Reg}$ is used to avoid collapse during anchor learning. Representations $g_{i}^{t}$ and $g_{i}^{s}$ are hence grounded into the learnable source anchors $\mathcal{A}_{s}$ through a novel Cross-Domain Attention layer taking visual representations as queries, source/target anchors as keys and source anchors as values. Classification is finally achieved through an MLP classifier supervised through a cross-entropy loss $\mathcal{L}_{CE}$ using ground truth labels in the source domain and pseudo-labels in the target domain. The visual encoder, Cross-Domain Attention, and MLP layers share weights across domains.
@@ 3 Method
In Stage 1, a language-based reference structure space is defined. We represent this structure as a set of anchors, which we will use as a reference to guide representation learning. In Stage 2, a text-supervisor language model is trained to map image/video captions to class categories in order to provide pseudo-labels for the unlabeled target domain. The language model is also trained to keep the internal text embeddings aligned to the domain-agnostic structure defined in Stage 1. Finally, in Stage 3, a cross-domain visual classifier is trained to categorize source and target samples by learning domain-specific action anchors bound to be structurally similar to the domain-agnostic structure defined in Stage 1.
@@ 3.3 Stage 2: Training of the Language Supervisor
Next, we encourage the vector representation $z_{i}^{s}$ to be geometrically aligned to the anchor $\mathcal{A}[y_{i}^{s}]$ corresponding to ground truth action $y_{i}^{s}$. To do so, we compute $\mathbf{r}^{z_{i}^{s}}=rel(z_{i}^{s},\mathcal{A})$, the relative encoding of $z_{i}^{s}$, and supervise it to be similar to $\mathbf{r}^{y_{i}^{s}}=rel(\mathcal{A}[y_{i}^{s}],\mathcal{A})$, the relative encoding of the anchor $\mathcal{A}[y_{i}^{s}]$, with the following structure-preserving loss:
@@ 3.4 Stage 3: Cross-Domain Visual Classifier
Stage 3 aims to train a cross-domain visual classifier aligning representations extracted through a visual encoder $\mathcal{V}$ to the structure imposed by the domain-agnostic language anchors $\mathcal{A}$. We allow each domain to develop its own latent space, but we encourage both spaces to be aligned to two sets of learnable anchors $\mathcal{A}_{t}\in\mathbb{R}^{N_{c}\times D_{v}}$ and $\mathcal{A}_{s}\in\mathbb{R}^{N_{c}\times D_{v}}$ specific to the target and source domain respectively. We further ensure that the structures of $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$ are aligned to the domain-agnostic, fixed anchors $\mathcal{A}$ through the supervision provided by ground truth labels $y_{i}^{s}$ in the source domain and the text supervision provided by $\mathcal{G}$ in the target domain in the form of the predicted vector $z_{i}^{t}$ and the predicted pseudo-label $\overline{y}_{i}^{t}$. The source and target datasets are merged ($\hat{\mathcal{X}}=\mathcal{X}_{t}+\mathcal{X}_{s}$) and used for training with a total of $M=N_{s}+N_{t}$ samples. Source samples comprise an image, caption, and label triplet $(x_{i}^{s},l_{i}^{s},y_{i}^{s})$, whereas target samples $(x_{i}^{t},l_{i}^{t},\overline{y}_{i}^{t})$ replace label $y_{i}^{s}$ with the pseudo-label $\overline{y}_{i}^{t}$. Target and source images are processed by $\mathcal{V}$ to obtain latent representations $g_{i}^{t}$ and $g_{i}^{s}$, while target captions $l_{i}^{t}$ are processed by $\mathcal{G}$ to obtain latent representations $z_{i}^{t}$ as follows:
$g_{i}^{t}=\mathcal{V}(x_{i}^{t}),\hfill z_{i}^{t}=\mathcal{G}(l_{i}^{t}),\hfill g_{i}^{s}=\mathcal{V}(x_{i}^{s}).$ (5)
@@ 3.4.1 Structure Learning
This loss encourages the visual representations to maintain relationships with their respective domain-specific visual anchors that mirror the relationships expressed by the domain-invariant anchors $\mathcal{A}$. This process supervises both the learnable anchors $\mathcal{A}_{s},\mathcal{A}_{t}$ and the visual encoder $\mathcal{V}$. However, since this loss alone might lead to the collapse of anchor representations into a small region, hindering effective classification, we introduce a volume (or spread) regularization loss. This loss treats each set of anchors as a multidimensional parallelotope whose spread or volume occupied in the latent space can be measured by the determinant of its Gram matrix. Given the Gram matrices for the three sets of anchors:
$\gamma=\mathcal{A}\mathcal{A}^{T},\hfill\gamma_{s}=\mathcal{A}_{s}\mathcal{A}_{s}^{T},\hfill\gamma_{t}=\mathcal{A}_{t}\mathcal{A}_{t}^{T},$ (7)
we encourage the volume of each domain-specific parallelotope to be approximately equal to the volume of the reference anchors $\mathcal{A}$ with the following regularization loss: e
@@
where we consider log-determinants instead of determinants for numerical stability. This regularization loss ensures the visual anchors occupy a volume similar to the one of $\mathcal{A}$, preventing collapse and promoting better separation of classes in the representation space.
@@ 3.4.2 Classification Training
The network undergoes classification training concurrently with structure learning. To ensure that source and target visual representations are grounded in the structure imposed by $\mathcal{A}$, while still being free to capture domain-specific nuances, we propose a novel cross-domain attention layer that aims to ground domain-specific visual representations $g_{i}^{t}$ and $g_{i}^{s}$ to the structure of source anchors $\mathcal{A}_{s}$. We include two Cross-Domain attention layers sharing the same weights for the target and source branches of the architecture (See Fig. 2). In the target branch, we use $g_{i}^{t}$ as queries, target anchors $\mathcal{A}_{t}$ as keys, and source anchors $\mathcal{A}_{s}$ as values. The output is summed to $g_{i}^{t}$ with a residual connection to include both domain-specific and cross-domain information in the final representation:
@@
Crucially, the cross-domain attention layer constructs an attention map by leveraging the domain-specific relations between the encoded representations and the anchors (i.e., $QK^{T}$). These relations are expected to be similar across domains due to the influence of the structural loss $\mathcal{L}_{S}$. This process results in a unified representation since the output of the attention layer always depends on values coming from the source anchors $\mathcal{A}_{s}$. The residual connection included in the attention layer is introduced to account for domain-dependent class characteristics, which results in final feature vectors $f_{i}^{s}$ and $f_{i}^{t}$. Finally, an MLP classification head processes $f_{i}^{t}$ and $f_{i}^{s}$ to output predictions $\hat{y}_{i}^{t}$ and $\hat{y}_{i}^{s}$. The MLP has shared weights across domains and is optimized with a cross-entropy loss using the ground truth label $y_{i}^{s}$ for source examples and the pseudo-label $\overline{y}_{i}^{t}$ predicted by $\mathcal{G}$ in the case of target samples. The overall training objective of the classifier is the following:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{CE}(\hat{y_{i}},y_{i}^{*})+\lambda_{2}\mathcal{L}_{S}(\mathbf{r}^{*},\mathbf{r}^{g_{i}})+\lambda_{3}\mathcal{L}_{Reg}(\gamma,\gamma_{*}),$ (11)
@@ 4 Experiments
Table 1: Results on DomainNet [50], comprising $4$ domains: Real, Clipart, Sketch, and Painting, for a total of 12 domain adaptation scenarios. All models use the Swin-B [37] backbone. Best results are reported in bold, whereas second best results are underlined.
@@ 4.1 Experimental Setup
Datasets. We evaluate LAGUNA on four comprehensive UDA benchmarks: DomainNet [50], GeoPlaces [26] GeoImnet [26], and Ego2Exo [27, 20]. DomainNet consists of $400K$ images across $345$ classes and is used to evaluate performance in $12$ domain transfer settings [68, 27]. GeoImnet and GeoPlaces are subsets of the larger GeoNet dataset [26], contain over 750K images and focus on geographic disparities in image classification (GeoImNet with $600$ classes) and scene recognition (GeoPlaces with $205$ classes). Finally, Ego2Exo is a subset of EgoExo-4D curated for domain adaptation. It involves a total of $9086$ videos and $24$ action categories. DomainNet and GeoNet include per-image captions derived from metadata and BLIP-2 [32], respectively, while Ego2Exo provides video descriptions from the original EgoExo-4D dataset.
Model and training details. Our method comprises three base components: a language model for structure definition (Stage 1), a text supervisor model (Stage 2), and a visual backbone (Stage 3). We use a pre-trained SentenceTransformer [55] for structure definition and a BERT base model [59] as text supervisor. The latter is trained on the source domain of each scenario for 5 epochs with a batch size of 64, a learning rate of 1e-4, and the AdamW [41] optimizer. The visual encoders are tailored to each dataset. For DomainNet, we adapt the Swin-B backbone [37], following prior work [77, 73, 25]. For GeoImnet and GeoPLaces, we utilize the ViT-B backbone [12] following [68, 77]. For Ego2Exo, we employ pre-extracted Omnivore-L features [20] as in [27]. All vision models undergo joint training with the other network components for $10$ epochs with a batch size of $32$, an initial learning rate of $41e-$ 4, and cosine scheduling. We set the loss weights to $\lambda_{1}$=1.0 $\lambda_{2}$=0.1 and $\lambda_{3}$=0.001 to normalize the relative magnitude of the related loss functions.

### v3

@@ Abstract
Unsupervised domain adaptation remains a critical challenge in enabling the knowledge transfer of models across unseen domains. Existing methods struggle to balance the need for domain-invariant representations with preserving domain-specific features. This is often due to the fact that alignment approaches impose the projection of samples with similar semantics close to each other in the latent space, despite drastic domain differences. We introduce LAGUNA - LAnguage Guided UNsupervised Adaptation with structured spaces, a novel approach that shifts the focus from aligning representations in absolute coordinates to aligning the relative positioning of equivalent concepts in latent spaces. LAGUNA defines a domain-agnostic structure on the semantic/geometric relationships between class labels in the language space and guides the adaptation, ensuring that the organization of samples in visual space reflects reference inter-class relationships while preserving domain-specific characteristics. Remarkably, LAGUNA surpasses previous work in $18$ different adaptation scenarios across four diverse image and video datasets with average accuracy improvements of up to $+3.32\%$ on DomainNet, $+5.75\%$ on GeoPlaces, $+4.77\%$ on GeoImnet, and $+1.94\%$ mean class accuracy improvement on EgoExo4D.
@@ 1 Introduction
Based on this observation, we introduce LAGUNA - LAnguage Guided UNsupervised Adaptation with structured spaces, a novel approach to domain adaptation which guides source and target spaces to develop semantic-geometric inter-relationships reflecting the structure of a reference space (Figure 1(bottom)). As shown in recent works, language can provide a semantic space agnostic to the nuances of visual observations enabling robust zero-shot generalization [56, 10] and supporting domain robustness [13, 68, 23, 47, 27], hence we choose language to build our reference structured space in LAGUNA. This assumes the availability for both source and target samples of natural language descriptions, which can be generated from available captioning models [33] or collected from the web at a fraction of the cost of typical human labeling procedures [27]. LAGUNA employs a 3-stage approach to structurally align source and target representation spaces while allowing domains to preserve typical patterns (Figure 2). In Stage 1, textual class labels are mapped to a domain-agnostic reference space representing their relationships at the conceptual level. In Stage 2, a language model is trained to map textual captions to the reference latent space, providing pseudo-labels for target samples. In Stage 3, a cross-domain classifier is trained to encourage domain-specific representations to follow the reference structure.
Experiments demonstrate the superiority of LAGUNA over existing state-of-the-art methods across $18$ different domain adaptation scenarios sourced from four diverse image and video datasets, with gains of up to $+3.32\%$ on DomanNet [52], $+5.75\%$ on GeoPlaces [26], $+4.77\%$ on GeoImnet [26] and $+1.94\%$ mean per class accuracy on EgoExo4D [20]. We further report ablations to analyze the specific contributions of our design choices.
In sum, our main contributions are as follows: 1) we investigate the use of relative representations for UDA, showing its advantages with respect to the current absolute alignment paradigm; 2) we propose LAGUNA, a three-stage method which learns a cross-domain classifier where individual source and target spaces are distinct yet aligned to a reference structure derived from language; 3) through extensive ablations and comparisons with state-of-the-art models, we show the superiority of the proposed approach.
@@ 2.1 UDA in Computer Vision
Figure 2: The 3 stages of LAGUNA. In Stage 1, a domain-agnostic semantic structure in the form of a set of anchors $\mathcal{A}$ is defined. In Stage 2, a language model $\mathcal{G}$ is trained on source captions-label pairs to provide pseudo-labels for the target domain. The model is trained to be structurally aligned to anchors ${}^{*}\mathcal{A}$ with a structural loss $\mathcal{L}_{S}$ and to predict class labels with a cross-entropy loss $\mathcal{L}_{CE}$. In Stage 3, a cross-domain visual classifier is trained on labeled source and unlabeled target images. A visual encoder $\mathcal{V}$ is used to represent target and source samples as $g_{i}^{t}$ and $g_{i}^{s}$, respectively. The cosine similarities $\mathbf{r}^{g_{i}^{t}}$ between $g_{i}^{t}$ and learnable target anchors $\mathcal{A}_{t}$ are bound to be alike to the similarities $\mathbf{r}^{z_{i}^{t}}$ between textual representation of the target caption $z_{i}^{t}$ predicted via $\mathcal{G}$ and ${}^{*}\mathcal{A}$ by means of $\mathcal{L}_{S}$. A similar processing is applied to $g_{i}^{s}$, the textual representation of the ground truth class ${}^{*}A[y_{i}^{S}]$ and $\mathcal{A}_{s}$. A regularization loss $\mathcal{L}_{Reg}$ is used to avoid collapse during anchor learning. Representations $g_{i}^{t}$ and $g_{i}^{s}$ are hence grounded into the learnable $\mathcal{A}_{s}$ through a novel Cross-Domain Attention layer taking visual representations as queries, source/target anchors as keys and source anchors as values. Classification is finally achieved with an MLP classifier supervised via $\mathcal{L}_{CE}$ using ground truth labels in the source domain and pseudo-labels in the target domain.
@@ 3 Method
LAGUNA comprises a 3-stage process, as illustrated in Fig. 2. In Stage 1, a language-based reference structure space is defined. We represent this structure as a set of anchors, which we will use as a reference to guide representation learning. In Stage 2, a text-supervisor language model is trained to map image/video captions to class categories, providing pseudo-labels for the unlabeled target domain. The language model is also trained to keep internal text embeddings aligned to the domain-agnostic structure defined in Stage 1. In Stage 3, a cross-domain visual classifier is trained to categorize source and target samples by learning domain-specific action anchors that are structurally similar to the domain-agnostic structure defined in Stage 1.
@@ 3.3 Stage 2: Training of the Language Supervisor
Next, we encourage the vector representation $z_{i}^{s}$ to be geometrically aligned to $\mathcal{A}[y_{i}^{s}]$ corresponding to ground truth action $y_{i}^{s}$. To do so, we compute $\mathbf{r}^{z_{i}^{s}}=rel(z_{i}^{s},\mathcal{A})$, the relative encoding of $z_{i}^{s}$, and supervise it to be similar to $\mathbf{r}^{y_{i}^{s}}=rel(\mathcal{A}[y_{i}^{s}],\mathcal{A})$, the relative encoding of the anchor $\mathcal{A}[y_{i}^{s}]$, with the following structure-preserving loss:
@@ 3.4 Stage 3: Cross-Domain Visual Classifier
Stage 3 trains a cross-domain visual classifier aligning representations extracted through a visual encoder $\mathcal{V}$ to the structure imposed by the domain-agnostic language anchors $\mathcal{A}$. We allow each domain to develop its own latent space, but we encourage both spaces to be aligned to two sets of learnable anchors $\mathcal{A}_{t}\in\mathbb{R}^{N_{c}\times D_{v}}$ and $\mathcal{A}_{s}\in\mathbb{R}^{N_{c}\times D_{v}}$ specific to the target and source domain respectively. We further ensure that the structures of $\mathcal{A}_{t}$ and $\mathcal{A}_{s}$ are aligned to the structure of $\mathcal{A}$ through the supervision provided by ground truth labels $y_{i}^{s}$ in the source domain and the text supervision provided by $\mathcal{G}$ in the target domain in the form of the predicted vector $z_{i}^{t}$ and the predicted pseudo-label $\overline{y}_{i}^{t}$. The source and target datasets are merged ($\hat{\mathcal{X}}=\mathcal{X}_{t}+\mathcal{X}_{s}$) and used for training with a total of $M=N_{s}+N_{t}$ samples. Source samples comprise an image, caption, and label triplet $(x_{i}^{s},l_{i}^{s},y_{i}^{s})$, whereas target samples $(x_{i}^{t},l_{i}^{t},\overline{y}_{i}^{t})$ replace label $y_{i}^{s}$ with the pseudo-label $\overline{y}_{i}^{t}$. Target and source images are processed by $\mathcal{V}$ to obtain latent representations $g_{i}^{t}$ and $g_{i}^{s}$, while target captions $l_{i}^{t}$ are processed by $\mathcal{G}$ to obtain latent representations $z_{i}^{t}$ as follows:
$g_{i}^{t}=\mathcal{V}(x_{i}^{t}),\hskip 14.22636ptz_{i}^{t}=\mathcal{G}(l_{i}^{t}),\hskip 14.22636ptg_{i}^{s}=\mathcal{V}(x_{i}^{s}).$ (5)
@@ 3.4.1 Structure Learning
This loss encourages visual representations to maintain relationships with their respective domain-specific visual anchors that mirror the relationships expressed by the domain-invariant anchors $\mathcal{A}$. This process supervises both the learnable anchors $\mathcal{A}_{s},\mathcal{A}_{t}$ and the visual encoder $\mathcal{V}$. However, since this loss alone might lead to the collapse of anchor representations into a small region, hindering effective classification, we introduce a volume (or spread) regularization loss. This loss treats each set of anchors as a multidimensional parallelotope whose volume in the latent space can be measured by the determinant of its Gram matrix. Given the Gram matrices for the three sets of anchors:
$\gamma=\mathcal{A}\mathcal{A}^{T},\hskip 14.22636pt\gamma_{s}=\mathcal{A}_{s}\mathcal{A}_{s}^{T},\hskip 14.22636pt\gamma_{t}=\mathcal{A}_{t}\mathcal{A}_{t}^{T},$ (7)
we encourage the volume of each domain-specific parallelotope to be approximately equal to the volume of the reference anchors $\mathcal{A}$ with the following regularization loss:
@@
where we use log-determinants for numerical stability. This regularization loss ensures taht visual anchors occupy a volume similar to the one of $\mathcal{A}$, preventing collapse and promoting better class separation in the representation space.
@@ 3.4.2 Classification Training
We perform classification training concurrently with structure learning. To ensure that source and target visual representations are grounded in the structure imposed by $\mathcal{A}$, while still being free to capture domain-specific nuances, we propose a novel cross-domain attention layer that aims to ground domain-specific visual representations $g_{i}^{t}$ and $g_{i}^{s}$ to the structure of source anchors $\mathcal{A}_{s}$. We include two Cross-Domain attention layers sharing the same weights for the target and source branches of LAGUNA (See Fig. 2). In the target branch, we use $g_{i}^{t}$ as queries, target anchors $\mathcal{A}_{t}$ as keys, and source anchors $\mathcal{A}_{s}$ as values. The output is summed to $g_{i}^{t}$ with a residual connection to include both domain-specific and cross-domain information in the final representation:
@@
Crucially, the cross-domain attention layer constructs an attention map by leveraging the domain-specific relations between the encoded representations and the anchors (i.e., $QK^{T}$). These relations are expected to be similar across domains due to the influence of the structural loss $\mathcal{L}_{S}$. This process results in a unified representation since the output of the attention layer always depends on values coming from the source anchors $\mathcal{A}_{s}$. The residual connection included in the attention layer is introduced to account for domain-dependent class characteristics, which results in the feature vectors $f_{i}^{s}$ and $f_{i}^{t}$. Finally, an MLP classification head processes $f_{i}^{t}$ and $f_{i}^{s}$ to output predictions $\hat{y}_{i}^{t}$ and $\hat{y}_{i}^{s}$. The MLP has shared weights across domains and is optimized with a cross-entropy loss using the ground truth label $y_{i}^{s}$ for source examples and the pseudo-label $\overline{y}_{i}^{t}$ predicted by $\mathcal{G}$ in the case of target samples. The overall training objective is:
$\small\mathcal{L}=\lambda_{1}\mathcal{L}_{CE}(\hat{y_{i}},y_{i}^{*})+\lambda_{2}\mathcal{L}_{S}(\mathbf{r}^{*},\mathbf{r}^{g_{i}})+\lambda_{3}\mathcal{L}_{Reg}(\gamma,\gamma_{*}),$ (11)
@@ 4 Experiments
ViT-B CLIP* [56] 72.39 60.90 66.81 81.37 60.90 66.81 81.37 72.39 66.81 81.37 72.39 69.90 70.38
DAPrompt* [16] 73.90 65.90 70.40 84.90 65.80 70.20 84.60 73.50 69.90 84.90 73.80 65.80 73.80
PADCLIP [31] 76.40 67.50 72.70 84.20 68.10 71.10 83.60 76.30 71.70 83.50 75.40 67.20 74.81
[…]
UniMoS* [36] 77.50 68.00 72.50 86.00 68.50 72.30 85.90 77.80 72.60 85.80 77.20 68.20 76.03
LAGUNA 79.53 73.18 74.24 86.04 73.54 74.98 85.81 80.24 74.43 86.57 79.89 74.11 78.54
Improvement +2.03 +4.07 +1.74 +0.04 +5.03 +2.68 -0.09 +2.44 +1.83 +0.77 +2.69 +5.91 +2.51
Table 1: Results on DomainNet [52] in $4$ domains: Real, Clipart, Sketch, and Painting, for 12 adaptation scenarios using the Swin-B [39] and ViT-B [12] backbones. The best results are in bold, second bests are underlined.‘*‘ method uses domain-aware prompts, $\dagger$ our run.
@@ 4.1 Experimental Setup
Datasets. We evaluate LAGUNA on four comprehensive UDA benchmarks: DomainNet [52], GeoPlaces [26] GeoImnet [26], and Ego2Exo [27, 20]. DomainNet consists of $400K$ images across $345$ classes and is used to evaluate performance in $12$ domain transfer settings [70, 27]. GeoImnet and GeoPlaces are subsets of the larger GeoNet dataset [26], contain over 750K images and focus on geographic disparities in image classification (GeoImNet with $600$ classes) and scene recognition (GeoPlaces with $205$ classes). Finally, Ego2Exo is a subset of EgoExo-4D [20] curated for domain adaptation. It involves a total of $9086$ videos and $24$ action categories. DomainNet and GeoNet include per-image captions derived from BLIP-2 [33] and web metadata, respectively, while Ego2Exo provides video descriptions from the original EgoExo-4D dataset.
Model and training details. We use a pre-trained SentenceTransformer [57] for structure definition (Stage 1) and a BERT base model [61] as text supervisor (Stage 2). The latter is trained on the source domain of each scenario for 5 epochs with a batch size of 64, a lr of 1e-4, and the AdamW [43] optimizer. The visual encoders (Stage 3) are tailored to each dataset for fair comparisons with existing approaches. For DomainNet, we adapt the Swin-B and ViT-B backbones [39, 12]. For GeoImnet and GeoPLaces, we utilize the ViT-B backbone [12] following [70, 79]. For Ego2Exo, we employ pre-extracted Omnivore-L features [20] as in [27]. All vision models undergo joint training with the other network components for $10$ epochs with a batch size of $32$, an initial learning rate of $41e-$ 4, and cosine scheduling. We set the loss weights to $\lambda_{1}$=1.0 $\lambda_{2}$=0.1 and $\lambda_{3}$=0.001 to normalize the relative magnitude of the related loss functions.
@@
4.2 Results
We benchmark our model against various language-free domain adaptation methods that have reported results on the considered datasets [79, 70, 25, 77, 7, 75]. Besides classic UDA methods, we also compare with previous approaches using language guidance [27, 56, 16, 31, 36]. Following the evaluation protocols of each benchmark, we report accuracy for DomainNet, GeoImNet, and GeoPlaces, and per-class mean accuracy for Ego2Exo.
DomainNet. Table 1 presents the results on the DomainNet dataset, encompassing $12$ domain adaptation scenarios across four distinct domains: Real (R), Clipart (C), Sketch (S), and Painting (P). LAGUNA is compared against previous state-of-the-art methods utilizing the Swin-B and ViT-B backbones as originally adopted by the compared methods to ensure fair comparison. LAGUNA significantly outperforms previous methods in 12 scenarios, with an average performance increase of $3.32\%$ for the Swin-B backbone. For ViT-B, LAGUNA leads in 11 of 12 scenarios, boosting performance by an average of $2.51\%$. Importantly, LAGUNA achieves SOTA ViT-B performance without relying on extensive CLIP pretraining and domain-specific prompting, unlike previous models [31, 36, 16].
[…]
These consistent improvements across all tested datasets, with a total of 18 scenarios, underscore the effectiveness of LAGUNA with respect to both language-free and language-guided UDA approaches.